LASER INDUCED FLUORESENCE SPECTROSCOPY OF THE JET-COOLED TRANSIENT SPECIES AsD_{2} AND AsHD

\underline{ROBERT A. GRIMMINGER} and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY, 40506-0055.

The Ã^{2}A_{1} − X̃^{2}B_{1} electronic band system of the AsD_{2} and AsHD free radicals has been observed by laser induced fluorescence (LIF) spectroscopy. AsD_{2} (or AsHD) was produced in a pulsed supersonic expansion by electric discharge dissociation of a mixture of 5% AsD_{3} (or AsH_{x}D_{3−x} mixture) seeded in argon. Single vibronic level emission spectra have been obtained for both AsD_{2} and AsHD. The observed LIF transitions of AsD_{2} in the 411 - 505 nm region exhibit a long progression in the excited state bending vibration. High-resolution rotationally resolved spectra of the 0_{0}^{0} band of AsD_{2} were also recorded. This c-type band shows prominent ΔK_{a} = ±3 transitions due to the large change in the asymmetry parameter upon electronic excitation. Extensive hyperfine splittings were resolved which allowed the determination of the excited state Fermi contact parameter. The data have been used to refine the molecular structure of the AsH_{2} radical.